A market authority charges fees for parking according to the following schedule; N5 for bicycles; N10 for motor cycles and N20 for cars. For three days, the records show the following parked:

Bicycles Motor cycles Cars

Monday   150   180 500

Tuesday   90   105 400

Wednesday   102   130 450

Required: Arrange this information in matrix form and calculate the collections made for bicycles, motorcycles and cars, hence the total revenue collected for three days.


1
Expert's answer
2021-06-28T23:21:02-0400

"\begin{pmatrix}
 5 & 10 &20
\end{pmatrix} \times \begin{pmatrix}
 150 & 90 & 102 \\
 180 & 105 &130\\
 500 & 400 & 450
\end{pmatrix}=\begin{pmatrix}
 12550 & 9500 & 10800
\end{pmatrix}"



"12550+9500+10800=32850"